OpenStudy (aivantettet26):

An elevator having a weight of 5 tons accelerates upward at a rate of 5 ft/s^2. What is the tension on the cable if it is accelerating downward the same rate?

OpenStudy (amistre64):

mg - a id assume

OpenStudy (amistre64):

letting g = 32ft/s^2; and 2000 lbs per ton ...

OpenStudy (aivantettet26):

7620 N = T

OpenStudy (aivantettet26):

T =F = ma T = 5000 kg x 1.524 m/s^2 T = 7620 N

OpenStudy (ybarrap):

If it is accelerating downward at a rate of 5 ft/s^2 then since it is not accelerating downward at a rate of 32.2 ft/s^2 (i.e. acceleration of gravity/free-fall), then there will be a force felt and it will be dependent on the difference between the force it would be at rest, mg, and the force it is experiencing, m*5; so, anyone in the elevator would feel lighter and the tension in the cable would also be lower than if the elevator were not moving (i.e. at rest): mg - m*5 = m(32.2 - 5) lbs*ft/s^2= [(2*5*1000)/32.2]*(32.2-5) ft lbs